Please help..
Solve the equation:
4^2x+1 = 2^3x+6

Answer by rmromero(383)


Solve the equation:
4^2x+1 = 2^3x+6
2^2(2x+1) = 2^(3x+6)
2^(4x+2) = 2^(3x+6) Notice that they have the same base 2
4x + 2 = 3x + 6 exponent = exponent
4x - 3x = 6 - 2 Solve or x
x = 4

Checking:
4^(((2)(4))+1) = 2^(((3)(4))+6)
4^9 = 2^18
262144 = 262144 ----> True